# Spectral Adapter: Fine-Tuning in Spectral Space

### v1

@@ Abstract
Recent developments in Parameter-Efficient Fine-Tuning (PEFT) methods for pretrained deep neural networks have captured widespread interest. In this work, we study the enhancement of current PEFT methods by incorporating the spectral information of pretrained weight matrices into the fine-tuning procedure. We investigate two spectral adaptation mechanisms, namely additive tuning and orthogonal rotation of the top singular vectors, both are done via first carrying out Singular Value Decomposition (SVD) of pretrained weights and then fine-tuning the top spectral space. We provide a theoretical analysis of spectral fine-tuning and show that our approach improves the rank capacity of low-rank adapters given a fixed trainable parameter budget. We show through extensive experiments that the proposed fine-tuning model enables better parameter efficiency and tuning performance as well as benefits multi-adapter fusion. The code will be open-sourced for reproducibility.
@@ 4.1 Language Model Fine-Tuning: Enhancing Fine-Tuning Results with Spectral AdapterA
Pre-Trained -- $38.82$
LoRAr=8 $0.16\%$ $46.25$
DoRAr=8 $0.17\%$ $45.79$
@@
We take $r=8$ for LoRA, DoRA and Spectral AdapterA to maintain approximately same number of trainable parameters for all methods. Table 2 presents the accuracy comparison where SpectralA stands for Spectral AdapterA. From the result, we observe that our Spectral AdapterA scores higher than both LoRA and DoRA by a large margin and increases the pretrained model baseline significantly, which verifies the effectiveness of the proposed spectral adaptation mechanism. See Appendix F.4 for more about experimental details.
@@ 4.2 Diffusion Model Fusion: Improving Multi-Object Fine-Tuning with Spectral AdapterA
We follow default training setting in [12] and fine-tune the Chilloutmix diffusion model [8] on three custom animal concepts, see original animals in "reference" in Figure 5. For better spatial alignment, we adopt T2I-Adapter [39] with sketch condition and we set guidance equal to one, see also "reference" in Figure 5 for the sketch condition being used. LoRA rank $r=8$ is adopted. For baseline comparisons, we use original code for Gradient Fusion [12] and Orthogonal Adaptation [42]. We adapt code of Gradient Fusion for FedAvg method since there is no official implementation available. Custom animal name is replaced with special token <$V_{\text{animal}}$> for fine-tuning. For our Spectral AdapterA, we follow the method depicted in Figure 4 and tune first, second, and third top eighth columns of singular vector matrices for different animal concepts. Figure 5 shows the generation results with different methods for selected prompts. Notably, baseline methods sometimes fail to capture the custom animal concepts while Spectral AdapterA recognizes all custom animals and generates visually satisfactory images.
@@ 4.3 Diffusion Model Expressiveness: Improving Parameter Efficiency with Spectral AdapterR
We fine-tune the Chilloumix diffusion model [8] with various PEFT methods on custom vase concept and present the generation results for prompt "a <$V_{\text{vase}}$>" in Figure 6 for various trainable parameter budgets, where grey dash denotes that the corresponding parameter budget is unobtainable with a given adapter no matter how the hyperparameter is chosen and empty entry without grey dash represents that there is a way to achieve the corresponding parameter budget though the generation result is skipped for better visualization. We follow default LoRA implementation in [12] for LoRA baseline and adjust it for all other methods. From Figure 6, it can be observed that LoRA, OFT, and LiDB start to generate vase close to custom vase with around $400\sim 500k$ trainable parameters.
Figure 6: Generation results for prompt “a <$V_{\text{vase}}$>” after fine-tuning Chilloutmix diffusion model [8] on custom vase images with different PEFT methods. See Section 4.3 for details.
SVDiff and VeRA generate unsatisfactory vase images. On the contrary, Spectral AdapterR starts to recognize the custom vase concept with only $200\sim 300k$ trainable parameters and has finer-grained parameter choices compared to other methods, i.e., notably Spectral AdapterR can have as few as $1k$ parameters while other methods start with at least tens of thousands of trainable parameters. In a word, Spectral AdapterR enjoys finer-grained parameter budget choices and manifests better visual quality with fewer parameters, thus achieves enhanced parameter efficiency compared to various other PEFT methods. See Appendix F.6 for more generation results.
@@ F.1 Experimental Setup for Figure 1
MNLI $55e-$ 5 32 7 AdamW 0.01
RTE $42e-$ 4 32 15 AdamW 0.01
QNLI $56e-$ 5 32 7 AdamW 0.01
@@
QQP $41e-$ 4 32 8 AdamW 0.01
SST-2 $41e-$ 4 32 10 AdamW 0.01
CoLA $43e-$ 4 32 5 AdamW 0.01
STS-B $42e-$ 4 32 12 AdamW 0.01
@@
Table 4 shows the hyperparameter setting for our Spectral AdapterA used for fine-tuning DeBERTaV3-base model in Section 4.1. Rank $8$ is adopted for LoRA, DoRA, AdaLoRA and Spectral ${}^{A}.$ We set number of diagonal blocks to be $4$ and enable block sharing for OFT to maintain similar amount of trainable parameters.
@@ F.3 More About DeBERTaV3-base Experiment
LoRA $52.5e-$ 5 16 4 5 0.05 0.01
DoRA $52.5e-$ 5 16 4 5 0.05 0.01
Spectral AdapterA $52.5e-$ 5 - 4 5 - 0.01
@@
Table 5 shows training hyperparameter setting for fine-tuning Mistral 7B model in Section 4.1. We train with bfloat16 precision and fine-tune all q__\__proj, k__\__proj, v__\__proj, o__\__proj, and gate__\__proj weights. We evaluate with lm-evaluation-harness [46].
@@ F.5.2 More Multi-Adapter Fusion Generation Results
Method learning rate
LoRA $51e-$ 5
VeRA $35e-$ 3
OFTA $51e-$ 5
LiDB $45e-$ 4
SVDiff $31e-$ 3
Spectral AdapterR $43e-$ 4
Table 6: Hyperparameters for diffusion model fine-tuning task in Section 4.3
In this section, we present more tuning results with various parameter budgets for parameter efficiency experiment studied in Section 4.3, see Section 4.3 for baseline method explanation. Table 6 shows the learning rates used for each method, the rest experimental settings are default as in [12]. Figure 14 presents generation results of Chilloutmix diffusion model [8] tuned on custom chair concept with different methods under various parameter budgets. The prompt used is "a yellow <$V_{\text{chair}}$>". See "reference" in Figure 14 for original chair images. From the generation results, it can be observed that LoRA generates reasonable chairs for all rank $r=1,2,3$ though the generation quality is not stable, i.e., $r=3$ is worse than $r=2$. OFT and VeRA start to recognize custom chair with $>300k$ parameters, SVDiff generates unsatisfactory image and has fixed trainable parameters. LiDB doesn’t generate good results even for parameter budget $>800k.$ On the contrary, our Spectral AdapterR generates reasonable images with parameters as few as $\sim 100k$ and produces stable custom chairs since then.
Figure 15 shows generation results of Chilloutmix diffusion model [8] fine-tuned on custom table concept with different methods under various parameter budgets. The prompt used is “a <$V_{\text{table}}$>”. LoRA generates good images for all rank $r=1,2,3$ though it starts with $273k$ parameters even if rank is set to $1$. OFT also generates desirable images for all budgets though the fewest number of parameters it can achieve is $29k$. VeRA and LiDB start to generate reasonable images with $>300k$ trainable parameters and SVDiff doesn’t recognize the custom table accurately. Meanwhile, our Spectral AdapterR recognizes the shape of custom table with only $1k$ parameters and produces ideal images since $20k$ parameters.
@@
[…]

### v2

@@ Abstract
Recent developments in Parameter-Efficient Fine-Tuning (PEFT) methods for pretrained deep neural networks have captured widespread interest. In this work, we study the enhancement of current PEFT methods by incorporating the spectral information of pretrained weight matrices into the fine-tuning procedure. We investigate two spectral adaptation mechanisms, namely additive tuning and orthogonal rotation of the top singular vectors, both are done via first carrying out Singular Value Decomposition (SVD) of pretrained weights and then fine-tuning the top spectral space. We provide a theoretical analysis of spectral fine-tuning and show that our approach improves the rank capacity of low-rank adapters given a fixed trainable parameter budget. We show through extensive experiments that the proposed fine-tuning model enables better parameter efficiency and tuning performance as well as benefits multi-adapter fusion. Code is released at https://github.com/pilancilab/spectral_adapter.
@@ 4.1 Language Model Fine-Tuning: Enhancing Fine-Tuning Results with Spectral AdapterA
Pre-Trained -- $37.91\pm 1.34$
LoRAr=8 $0.16\%$ $44.81\pm 1.37$
DoRAr=8 $0.17\%$ $43.82\pm 1.37$
@@
We take $r=8$ for LoRA, DoRA and Spectral AdapterA to maintain approximately same number of trainable parameters for all methods. Table 2 presents the accuracy comparison where SpectralA stands for Spectral AdapterA. From the result, we observe that our Spectral AdapterA scores higher than both LoRA and DoRA by a large margin and increases the pretrained model baseline significantly, which verifies the effectiveness of the proposed spectral adaptation mechanism. See Appendix F.4 for more about experimental details. Note for a different learning rate, DoRA performs better than LoRA while still worse than our method, see also Appendix F.4 for details.
@@ 4.2 Diffusion Model Fusion: Improving Multi-Object Fine-Tuning with Spectral AdapterA
We follow default training setting in [12] and fine-tune the Chilloutmix diffusion model [8] on three custom animal concepts, see original animals in "reference" in Figure 5. For better spatial alignment, we adopt T2I-Adapter [39] with sketch condition and we set guidance equal to one, see also "reference" in Figure 5 for the sketch condition being used. LoRA rank $r=8$ is adopted. For baseline comparisons, we use original code for Gradient Fusion [12] and Orthogonal Adaptation [42]. We adapt code of Gradient Fusion for FedAvg method since there is no official implementation available. Custom animal name is replaced with special token <$V_{\text{animal}}$> for fine-tuning. For our Spectral AdapterA, we follow the method depicted in Figure 4 and tune first, second, and third top eighth columns of singular vector matrices for different animal concepts. Figure 5 shows the generation results with different methods for selected prompts. Notably, baseline methods sometimes fail to capture the custom animal concepts while Spectral AdapterA recognizes all custom animals and generates visually satisfactory images. For better measurement, we also compute the alignment scores for each generated image with both reference images and prompt texts. It can be witnessed that our method achieves better alignment scores compared to baselines. See Appendix F.7 for details on alignment score computation.
@@ 4.3 Diffusion Model Expressiveness: Improving Parameter Efficiency with Spectral AdapterR
Figure 6: Generation results for prompt “a <$V_{\text{vase}}$> on a table” after fine-tuning Chilloutmix diffusion model [8] on custom vase images with different PEFT methods. See Section 4.3 for details.
We fine-tune the Chilloumix diffusion model [8] with various PEFT methods on custom vase concept and present the generation results for prompt "a <$V_{\text{vase}}$> on a table" in Figure 6 for various trainable parameter budgets, where grey dash denotes that the corresponding parameter budget is unobtainable with a given adapter no matter how the hyperparameter is chosen and empty entry without grey dash represents that there is a way to achieve the corresponding parameter budget though the generation result is skipped for better visualization. We follow default LoRA implementation in [12] for LoRA baseline and adjust it for all other methods. From Figure 6, it can be observed that LoRA, OFT, and LiDB start to generate vase close to custom vase with at least $200k$ trainable parameters. SVDiff and VeRA are unable to generate ideal vase images even if scaled to large parameter budget. On the contrary, Spectral AdapterR starts to recognize the custom vase concept with only $20k$ trainable parameters and has finer-grained parameter choices compared to other methods, i.e., notably Spectral AdapterR can have as few as $1k$ parameters while other methods start with at least tens of thousands of trainable parameters. In a word, Spectral AdapterR enjoys finer-grained parameter budget choices and manifests better visual quality with fewer parameters, thus achieves enhanced parameter efficiency compared to various other PEFT methods.
[…]
Figure 7 above presents generation results of Chilloutmix diffusion model [8] tuned on custom chair concept with different methods under various parameter budgets. The prompt used is "a yellow <$V_{\text{chair}}$>". See "reference" in Figure 7 for original chair images. From the generation results, it can be observed that LoRA generates reasonable chairs for all rank $r=1,2,3$ though it already induces $273k$ parameters even if rank is set to $1$. OFT and VeRA start to recognize custom chair with $>100k$ parameters. SVDiff has a single fixed trainable parameter budget of size around $100k$. LiDB forms a competitive candidate and generates satisfactory images with smallest trainable parameter budget among all baseline methods. However, our Spectral AdapterR still generates images better aligned to reference images with as few as $20k$ trainable parameters and has finer-grained parameter budget choices compared to LiDB. See Appendix F.6 for hyperparameter setting and Appendix F.7 for alignment score computation details.
@@ 5 Conclusion and Limitations
A limitation of the current work remains in the choice of tuning top spectral space. Though its validity has been theoretically verified under simple settings, further investigation on tuning different columns of singular vector matrices is critical to understanding the role of spectral information in fine-tuning procedure. Besides, fine-tuning spectral representation of different components, i.e., only the attention layer, of current large models is also worth studying. Moreover, the time consumption of singular value decomposition procedure increases as model grows larger and thus faster singular value decomposition method also benefits.
@@ F.1 Experimental Setup for Figure 1
MNLI $41e-$ 4 32 1 AdamW 0.01
RTE $43e-$ 4 32 10 AdamW 0.01
QNLI $41e-$ 4 32 1 AdamW 0.01
@@
QQP $41e-$ 4 32 10 AdamW 0.01
SST-2 $41e-$ 4 32 5 AdamW 0.01
CoLA $43e-$ 4 32 8 AdamW 0.01
STS-B $45e-$ 4 32 30 AdamW 0.01
@@
Table 4 shows the hyperparameter setting for our Spectral AdapterA used for fine-tuning DeBERTaV3-base model in Section 4.1. We set number of diagonal blocks to be $4$ and enable block sharing for OFT to maintain similar amount of trainable parameters.
@@ F.3 More About DeBERTaV3-base Experiment
LoRA $52.5e-$ 5 16 4 2 0.05 0.01
DoRA $52.5e-$ 5 16 4 2 0.05 0.01
Spectral AdapterA $52.5e-$ 5 - 4 2 - 0.01
@@
Table 5 shows training hyperparameter setting for fine-tuning Mistral 7B model in Section 4.1. We train with bfloat16 precision and fine-tune all q__\__proj, k__\__proj, v__\__proj, o__\__proj, and gate__\__proj weights. We evaluate with lm-evaluation-harness [47]. Table 6 shows accuracy comparison of different tuning methods with learning rate $51e-$ 5 . Our Spectral AdapterA still exceeds both LoRA and DoRA.
Method ##\##Param GSM8K
Pre-Trained -- $38.82$
LoRAr=8 $0.16\%$ $43.29\pm 1.36$
DoRAr=8 $0.17\%$ $43.52\pm 1.37$
[…]
Table 6: Supplemental experiments of fine-tuning Mistral 7B model with different PEFT methods with a different learning rate on GSM8K benchmark. See Section F.4 for experimental details.
@@ F.5.2 More Multi-Adapter Fusion Generation Results
In this section, we present more tuning results with various parameter budgets for parameter efficiency experiment studied in Section 4.3, see Section 4.3 for baseline method explanation. Table 7 shows the learning rates used for each baseline method and Table 8 shows learning rates used for our method, the rest experimental settings are default as in [12].
Method text encoder lr unet lr
LoRA $51e-$ 5 $41e-$ 4
VeRA ($r=1$) $31e-$ 3 $41e-$ 4
VeRA ($r=1024,4096$) $35e-$ 3 $41e-$ 4
OFTA $51e-$ 5 $41e-$ 4
LiDB $45e-$ 4 $41e-$ 4
SVDiff $31e-$ 3 $41e-$ 4
Table 7: Hyperparameters for baseline methods for diffusion model fine-tuning task in Section 4.3
Method vase chair table
text unet text unet text unet
Spectral AdapterR ($r=2,40$) $31e-$ 3 $21e-$ 2 $21e-$ 2 $21e-$ 2 $31e-$ 3 $21e-$ 2
Spectral AdapterR ($r=4$) $35e-$ 3 $35e-$ 3 $31e-$ 3 $21e-$ 2
Spectral AdapterR ($r=8$) $45e-$ 4 $25e-$ 2 $31e-$ 3 $21e-$ 2 $31e-$ 3 $21e-$ 2
Spectral AdapterR ($r=16$) $21e-$ 2 $31e-$ 3 $31e-$ 3 $21e-$ 2
Spectral AdapterR ($r=24$) $41e-$ 4 $21e-$ 2 $31e-$ 3 $31e-$ 3 $41e-$ 4 $21e-$ 2
Spectral AdapterR ($r=32$) $41e-$ 4 $25e-$ 2
Table 8: Hyperparameters for Spectral AdapterR for diffusion model fine-tuning task in Section 4.3
Figure 15 shows generation results of Chilloutmix diffusion model [8] fine-tuned on custom table concept with different methods under various parameter budgets. The prompt used is “a <$V_{\text{table}}$>”. LoRA generates acceptable images for all rank $r=1,2,3$ though it starts with $273k$ parameters even if rank is set to $1$. OFT generates desirable images only for parameter budget $>400k$. VeRA and LiDB start to generate reasonable images with $>300k$ trainable parameters and SVDiff has only a single fixed parameter budget. Meanwhile, our Spectral AdapterR recognizes the shape of custom table with as few as $6k$ parameters and produces ideal images since $100k$ parameters. See Appendix F.7 for alignment score computation details.
@@
F.7 Alignment Score Computation
For better quantitative measurement, we compute alignment scores for our Figure 5,6,7,15 results. Specifically, we first compute CLIP [46] embedding for all generated/reference images and prompt texts, then we compute the cosine similarity between generated images’ embedding and reference images’ embedding to serve as their alignment score. Likewise, text score stands for cosine similarity between generated images’ embeddings and their corresponding prompt texts’ embeddings. Intuition here is that if an image is close to another image (or text), their CLIP vectors are expected to stay close as well. For Figure 5 alignment score computation, we crop each generated image vertically into three columns, then we compute their alignment scores to each corresponding reference animal, we finally take the mean of these three scores. For Figure 6, 7, 15 scores, we compute average score over three random trials, with each trial consisting of $8$ generated images.